# Out-of-time ordered correlation functions for the localized $f$ electrons in the Falicov-Kimball model

### v1

@@ Abstract
We provide an exact evaluation of the out-of-time correlation (OTOC) functions for the localized $f$-particle states in the Falicov-Kimball model within dynamical mean-field theory. Different regimes of quantum chaos and quantum scrambling are distinguished by the winding numbers of the block Toeplitz matrices used in the calculation. The similarities of these fermionic OTOCs and their Lyapunov exponents for time evolution with the OTOCs for quantum spin models with disorder are also discussed.
@@ I Introduction
For chaotic systems, the initially local operator $W_{0}(t)$ expands, under Heisenberg dynamics, with a speed called the butterfly velocity and overlaps later with the operator $V_{r}$ resulting from the growth of the OTOC function inside the emergent light-cone in space-time. The growth rate of the OTOC is influenced by the interplay between quantum fluctuations and local scrambling, and depends on the size of the system and on its temperature. For thermal quantum systems with a large number of degrees of freedom, the exponential growth rate is characterized by the Lyapunov exponent bounded by the temperature $\lambda_{L}\leq 2\pi k_{B}T/\hbar$ —this bound is a bound on the growth rate of chaos [8, 9].
@@
which is equal to the Lyapunov exponent (obeying the chaos bound for “fast scramblers”) and can provide detailed behavior of the OTOC at its wavefront. Relaxation dynamics in many-body systems can be characterized by different time scales $\tau$, each with their own temperature and wave-vector dependencies, which affect the scrambling, and result in an anomalous behavior for the Lyapunov exponents [14].
@@
In this article, we investigate OTOC functions constructed from $f$-electron creation and annihilation operators (state-flip operators for the two-level system) for the FKM on an infinite dimension hypercubic lattice. First, we introduce the model and recall the derivation of the $f$-electron DOS. Next, we consider the lattice gas model and obtain analytic results for the single-particle DOS and OTOCs for any lattice, and then examine the infinite dimensional limit. Finally, we calculate OTOCs for $f$-electrons for the FKM and provide numerical results.
@@ II Model Hamiltonian and single-particle propagators
At half filling and for $U<0.866$, when both $1-U\operatorname{Re}\mathcal{G}_{0}^{t}(\omega=0)>0$ and $1+U\operatorname{Re}\mathcal{G}_{1}^{t}(\omega=0)>0$, the winding number is equal to zero (see Appendix A). Its logarithm is well defined and the corresponding determinants decay exponentially at large times for high temperatures, while they behave as a power-law for low temperatures (with a singular power law at $T=0$ only). This behavior can be seen in the change of the asymptotics of the logarithmic derivative from a constant to a $c/t$ dependence (Fig. 1). The corresponding DOS displays a Lorentzian peak for high temperatures, which is replaced by a power law singularity in the limit as $T\to 0$.
@@
where $W=\Delta t\to 0$ is the time discretization, and the long time asymptotic becomes a more complicated nonexponential one with oscillations of the logarithmic derivatives (Fig. 2). Such oscillations are strongest at high temperatures and decrease their amplitude and increase their period as the temperature is lowered.
@@ III Lattice gas
$G_{i}^{>}(t,0)$ $=-\mathrm{i}\left\langle U_{i}(t,0)c_{i}(0)c_{i}^{\dagger}(0)\right\rangle$
@@ IV.1 Anticommutator squared magnitude
$F_{ij}^{\textrm{mod}}(t)=\left\langle\left|[f_{i}(t),f_{j}^{\dagger}(0)]_{+}\right|^{2}\right\rangle$
$=\left\langle[f_{i}(t),f_{j}^{\dagger}(0)]_{+}[f_{i}(t),f_{j}^{\dagger}(0)]_{+}^{\dagger}\right\rangle$
$=\left\langle f_{i}(t)f_{j}^{\dagger}(0)f_{j}(0)f_{i}^{\dagger}(t)\right\rangle+\left\langle f_{j}^{\dagger}(0)f_{i}(t)f_{i}^{\dagger}(t)f_{j}(0)\right\rangle$
$+\left\langle f_{i}(t)f_{j}^{\dagger}(0)f_{i}^{\dagger}(t)f_{j}(0)\right\rangle+\left\langle f_{j}^{\dagger}(0)f_{i}(t)f_{j}(0)f_{i}^{\dagger}(t)\right\rangle$
$=\left\langle f_{i}f_{j}^{\dagger}f_{j}f_{i}^{\dagger}\right\rangle+\left\langle f_{j}^{\dagger}f_{i}f_{i}^{\dagger}f_{j}\right\rangle$
$+2\operatorname{Re}\left\langle f_{i}(t)f_{j}^{\dagger}(0)f_{i}^{\dagger}(t)f_{j}(0)\right\rangle.$ (45)
@@
$F_{ij}^{\textrm{mod}}(t)$ $=\left\langle f_{i}f_{j}^{\dagger}f_{j}f_{i}^{\dagger}\right\rangle+\left\langle f_{j}^{\dagger}f_{i}f_{i}^{\dagger}f_{j}\right\rangle$
$+2\operatorname{Re}\left\langle\mathcal{U}_{i}(t,0)f_{i}f_{j}^{\dagger}f_{i}^{\dagger}\mathcal{U}_{i}^{\dagger}(t,0)f_{j}\right\rangle,$ (46)
@@
$F_{ij}^{\textrm{mod}}(t)$ $=2\left\langle P_{i}^{-}P_{j}^{+}\right\rangle$
$+2\operatorname{Re}\left\langle\mathcal{U}_{i}(t,0)f_{i}f_{j}^{\dagger}f_{i}^{\dagger}\mathcal{U}_{i}^{\dagger}(t,0)f_{j}\right\rangle$
$=2\left\langle P_{i}^{-}P_{j}^{+}\right\rangle$
$-2\operatorname{Re}\left\langle P_{i}^{-}\mathcal{U}_{i}(t,0)f_{j}^{\dagger}\mathcal{U}_{i}^{\dagger}(t,0)f_{j}\right\rangle.$ (48)
@@ IV.2 Other trivial anticommutator functions
$F_{ij}^{\textrm{sq},c}(t)=\left\langle[f_{i}^{\dagger}(t),f_{j}^{\dagger}(0)]_{+}[f_{i}^{\dagger}(t),f_{j}^{\dagger}(0)]_{+}\right\rangle=0,$
$F_{ij}^{\textrm{sq},a}(t)=\left\langle[f_{i}(t),f_{j}(0)]_{+}[f_{i}(t),f_{j}(0)]_{+}\right\rangle=0.$ (50)
@@
$F_{ij}^{\textrm{mod},ca}(t)=\left\langle[f_{i}(t),f_{j}(0)]_{+}[f_{i}(t),f_{j}(0)]_{+}^{\dagger}\right\rangle$
$=\left\langle[f_{i}(t),f_{j}(0)]_{+}[f_{i}^{\dagger}(t),f_{j}^{\dagger}(0)]_{+}\right\rangle$
$=(1-\delta_{ij})2\left\langle P_{i}^{-}P_{j}^{-}\right\rangle$
$+(1-\delta_{ij})2\operatorname{Re}\left\langle\mathcal{U}_{i}(t,0)f_{i}f_{j}f_{i}^{\dagger}\mathcal{U}_{i}^{\dagger}(t,0)f_{j}^{\dagger}\right\rangle=0.$ (51)
@@ V Square of the anticommutator
$F_{ij}^{\textrm{sq}}(t)=\left\langle\left([f_{i}(t),f_{j}^{\dagger}(0)]_{+}\right)^{2}\right\rangle$
$=\left\langle[f_{i}(t),f_{j}^{\dagger}(0)]_{+}[f_{i}(t),f_{j}^{\dagger}(0)]_{+}\right\rangle$
$=\left\langle f_{i}(t)f_{j}^{\dagger}(0)f_{i}(t)f_{j}^{\dagger}(0)\right\rangle+\left\langle f_{j}^{\dagger}(0)f_{i}(t)f_{j}^{\dagger}(0)f_{i}(t)\right\rangle$
$+\left\langle f_{i}(t)f_{j}^{\dagger}(0)f_{j}^{\dagger}(0)f_{i}(t)\right\rangle+\left\langle f_{j}^{\dagger}(0)f_{i}(t)f_{i}(t)f_{j}^{\dagger}(0)\right\rangle.$ (52)
@@
$F_{ij}^{\textrm{sq}}(t)$ $=\left\langle\mathcal{U}_{i}(t,0)f_{i}f_{j}^{\dagger}\mathcal{U}_{i}(t,0)f_{i}f_{j}^{\dagger}\right\rangle$
$+\left\langle\tilde{\mathcal{U}}_{i}(t,0)f_{i}^{\dagger}f_{j}\tilde{\mathcal{U}}_{i}(t,0)f_{i}^{\dagger}f_{j}\right\rangle^{*}.$ (53)
@@
$F_{ii}^{\textrm{sq}}(t)$ $=\left\langle\mathcal{U}_{i}(t,0)\mathcal{U}_{i}(t,0)P_{i}^{-}\right\rangle$
$+\left\langle\tilde{\mathcal{U}}_{i}(t,0)\tilde{\mathcal{U}}_{i}(t,0)P_{i}^{+}\right\rangle^{*}.$ (55)
@@ VI Results and Discussion
At half filling $n_{d}=n_{f}=0.5$, the contributions $F_{0}(t)$ and $F_{1}(t)$ in (58) are equal and the total OTOC $F(t)$ is real. But, the partial contributions from the sites with empty $f$ particle states, $F_{0}(t)$, and with occupied $f$ particle states, $F_{1}(t)$, display intriguing time dependencies. For $U<0.44$, the real part of both characteristic functions (72) is positive at frequencies where the imaginary part vanishes and the winding number is zero (see Appendix B). The OTOC function (57) displays exponential decay at large times for high temperatures, which is replaced by a power law as $T\to 0$ (Fig. 4). The corresponding Lyapunov exponent (logarithmic derivative)
@@
approaches a constant value at large times for high temperatures within the chaos bound $|L_{0}(t)/(2\pi T)|<1$, whereas it displays $1/t$ dependence for low temperatures, which confirms the power law asymptotics of the OTOC. Consequently, the corresponding “density-of-states”
$A(\omega)=-\frac{1}{\pi}\operatorname{Im}\int_{0}^{+\infty}\mathrm{d}t\;\mathrm{e}^{\mathrm{i}\omega t}F_{ii}^{\textrm{sq}}(t)$ (74)
@@
Figure 4: (a) OTOC $F(t)$, (b,c) normalized Lyapunov exponent $L_{0}(t)/(2\pi T)$, and (d) the corresponding “density-of-states” $A(\omega)$ for $U=0.4$, $n_{f}=0.5$, $n_{d}=0.5$.
@@
where $W=\Delta t\to 0$ is the time discretization, and the long-time asymptotic becomes more complicated (Fig. 5) with oscillations of the logarithmic derivatives (Lyapunov exponent) $L_{0}(t)$ at high temperatures similar to the one for the single-particle propagators for $U>0.866$ (Fig. 2). With decreasing temperature, these oscillations are replaced by $1/t$ asymptotics, because we are still in the metallic phase $U<\sqrt{2}$. Now the OTOC’s “density-of-states” displays a two peak structure separated by a pseudogap at high temperatures with the development of the power law singularity in the center of the pseudogap as $T\to 0$. The sharp peak in the “density-of-states” is observed at low temperatures up to $U$ values where the “density-of-states” becomes negative, which reflects an increase in the quantum effects for these cases (such as Kirkwood-Dirac quasiprobabilities [36, 37, 38], which can also become negative).
For $U>0.866$, the imaginary part of the characteristic function (72) crosses the zero line at frequencies where $1\mp U\operatorname{Re}\mathcal{G}_{0,1}^{r}(\omega)=0$, but for $U<1.235$ the winding does not change because the real part is still positive at these points, whereas for $U>1.235$ the real part becomes negative and the winding number is equal to 2-2 (see Appendix B) resulting in a more complicated long time behavior of the OTOCs and their logarithmic derivatives (Lyapunov exponents) with damped oscillations (Fig. 6).
@@
It should be noted, that the Szegő-Widom theorem for block Toeplitz matrices [35] is applicable for smooth functions. For this case in the fermionic FKM, the functions are continuous at high temperatures. At low temperatures, both the matrix elements in (68) and characteristic functions (72) develop step-like features from the Fermi factors and the OTOC’s long time behavior is now governed by localization singularities given by the Anderson orthogonality catastrophe [39, 33, 32]. These cases of three different winding numbers for the characteristic function are also distinguished by the behavior of the Lyapunov exponent at low temperatures (see Appendix C, Fig. A.6).
At high temperatures, the oscillatory behavior of the Lyapunov exponent is similar to the one seen in [10, 11] for quantum spin models behind the light cone after the wave front passes. This analogy is plausible, because we consider a local OTOC function, which is always behind the wave front and at high temperature the effects of the Fermi statistics become negligible and the electrons behave like Pauli particles. Nevertheless, the localized $f$ particles act as impurity scattering centers for the itinerant $d$ electrons and the destruction of the oscillations of Lyapunov exponent for large Coulomb interaction $U$ is similar to the one observed in [11] with the an increase of disorder, when one observes the post wavefront ripples damped by disorder. This static disorder can protect the fast scrambling, which is manifested as a breaking of the bound on chaos, when at long times we have $L_{0}(t)>2\pi k_{B}T/\hbar$, which, is also manifested by the negative OTOC “density-of-states” $A(\omega)$, see Fig. 6. According to Yunger Halpern et al. [37], OTOCs can be considered as moments of the Kirkwood-Dirac quasiprobabilities which implies that negative values of the “density-of-states” $A(\omega)$ are signatures of quantum information scrambling.

### v2

@@ Abstract
We provide an exact evaluation of the out-of-time correlation (OTOC) functions for the localized $f$-particle states in the Falicov-Kimball model within dynamical mean-field theory. Different regimes of quantum chaos and quantum scrambling are distinguished by the winding numbers of the block Toeplitz matrices used in the calculation. The similarities of these fermionic OTOCs and their logarithmic derivatives for time evolution with the OTOCs for quantum spin models with disorder are also discussed.
@@ I Introduction
For chaotic systems, the initially local operator $W_{0}(t)$ expands, under Heisenberg dynamics, with a speed called the butterfly velocity and overlaps later with the operator $V_{r}$ resulting from the growth of the OTOC function inside the emergent light-cone in space-time. The growth rate of the OTOC is influenced by the interplay between quantum fluctuations and local scrambling, and depends on the size of the system and on its temperature. For thermal quantum systems with a large number of degrees of freedom, the initial classical exponential growth rate is characterized by the Lyapunov exponent bounded by the temperature $\lambda_{L}\leq 2\pi k_{B}T/\hbar$ —this bound is a bound on the growth rate of chaos [8, 9]. However, for quantum dissipative systems, e.g. systems in contact with a thermal environment, we find that instead of the short time exponential growth, they exhibit a long time decay with a decay rate that mirrors the classical Lyapunov exponent [10, 11].
@@
which is equal to the Lyapunov exponent for “fast scramblers” obeying the chaos bound, and can provide detailed behavior of the OTOC at its wavefront, even when it is not exponentially growing. Relaxation dynamics in many-body systems can be characterized by different time scales $\tau$, each with their own temperature and wave-vector dependencies, which affect the scrambling, and result in an anomalous behavior for the Lyapunov exponents [16].
@@
In this article, we investigate OTOC functions constructed from $f$-electron creation and annihilation operators (state-flip operators for the two-level system) for the FKM on an infinite dimension hypercubic lattice. First, we introduce the model and recall the derivation of the $f$-electron DOS. Next, we consider the lattice gas model and obtain analytic results for the single-particle DOS and OTOCs for any lattice, and then examine the infinite-dimensional limit. Finally, we calculate OTOCs for $f$-electrons for the FKM and provide numerical results.
@@ II Model Hamiltonian and single-particle propagators
At half filling and for $U<0.866$, when both $1-U\operatorname{Re}\mathcal{G}_{0}^{t}(\omega=0)>0$ and $1+U\operatorname{Re}\mathcal{G}_{1}^{t}(\omega=0)>0$, the winding number is equal to zero (see Appendix A). Its logarithm is well defined and the corresponding determinants decay exponentially at large times for high temperatures, while they behave as a power-law for low temperatures (with a singular power law at $T=0$ only). This behavior can be seen in the change of the asymptotics of the logarithmic derivative from a constant to a $c/t$ dependence (Fig. 1). The corresponding DOS displays a Lorentzian peak for high temperatures, which is replaced by a power law singularity in the limit as $T\to 0$ in agreement with the seminal results of Brandt and Urbanek [29].
@@
where $W=\Delta t\to 0$ is the time discretization, and the long time asymptotic becomes a more complicated nonexponential one with oscillations of the logarithmic derivatives (Fig. 2). Such oscillations are strongest at high temperatures and decrease their amplitude and increase their period as the temperature is lowered. Now, the corresponding DOS possesses a gap at low temperatures due to freezing of thermal fluctuations [29] and the presence of strong-correlations creating a Mott gap; the gap region is filled in by thermal excitations for increasing temperatures.
@@ III Lattice gas
$G_{i}^{>}(t,0)$ $=-\mathrm{i}\bigl\langle U_{i}(t,0)c_{i}(0)c_{i}^{\dagger}(0)\bigr\rangle$
@@ IV.1 Anticommutator squared magnitude
$F_{ij}^{\textrm{mod}}(t)=\bigl\langle\bigl|[f_{i}(t),f_{j}^{\dagger}(0)]_{+}\bigr|^{2}\bigr\rangle$
$=\bigl\langle[f_{i}(t),f_{j}^{\dagger}(0)]_{+}[f_{i}(t),f_{j}^{\dagger}(0)]_{+}^{\dagger}\bigr\rangle$
$=\bigl\langle f_{i}(t)f_{j}^{\dagger}(0)f_{j}(0)f_{i}^{\dagger}(t)\bigr\rangle+\bigl\langle f_{j}^{\dagger}(0)f_{i}(t)f_{i}^{\dagger}(t)f_{j}(0)\bigr\rangle$
$+\bigl\langle f_{i}(t)f_{j}^{\dagger}(0)f_{i}^{\dagger}(t)f_{j}(0)\bigr\rangle+\bigl\langle f_{j}^{\dagger}(0)f_{i}(t)f_{j}(0)f_{i}^{\dagger}(t)\bigr\rangle$
$=\bigl\langle f_{i}f_{j}^{\dagger}f_{j}f_{i}^{\dagger}\bigr\rangle+\bigl\langle f_{j}^{\dagger}f_{i}f_{i}^{\dagger}f_{j}\bigr\rangle$
$+2\operatorname{Re}\bigl\langle f_{i}(t)f_{j}^{\dagger}(0)f_{i}^{\dagger}(t)f_{j}(0)\bigr\rangle.$ (45)
@@
$F_{ij}^{\textrm{mod}}(t)$ $=\bigl\langle f_{i}f_{j}^{\dagger}f_{j}f_{i}^{\dagger}\bigr\rangle+\bigl\langle f_{j}^{\dagger}f_{i}f_{i}^{\dagger}f_{j}\bigr\rangle$
$+2\operatorname{Re}\bigl\langle\mathcal{U}_{i}(t,0)f_{i}f_{j}^{\dagger}f_{i}^{\dagger}\mathcal{U}_{i}^{\dagger}(t,0)f_{j}\bigr\rangle,$ (46)
@@
$F_{ij}^{\textrm{mod}}(t)$ $=2\bigl\langle P_{i}^{-}P_{j}^{+}\bigr\rangle$
$+2\operatorname{Re}\bigl\langle\mathcal{U}_{i}(t,0)f_{i}f_{j}^{\dagger}f_{i}^{\dagger}\mathcal{U}_{i}^{\dagger}(t,0)f_{j}\bigr\rangle$
$=2\bigl\langle P_{i}^{-}P_{j}^{+}\bigr\rangle$
$-2\operatorname{Re}\bigl\langle P_{i}^{-}\mathcal{U}_{i}(t,0)f_{j}^{\dagger}\mathcal{U}_{i}^{\dagger}(t,0)f_{j}\bigr\rangle.$ (48)
@@ IV.2 Other trivial anticommutator functions
$F_{ij}^{\textrm{sq},c}(t)=\bigl\langle[f_{i}^{\dagger}(t),f_{j}^{\dagger}(0)]_{+}[f_{i}^{\dagger}(t),f_{j}^{\dagger}(0)]_{+}\bigr\rangle=0,$
$F_{ij}^{\textrm{sq},a}(t)=\bigl\langle[f_{i}(t),f_{j}(0)]_{+}[f_{i}(t),f_{j}(0)]_{+}\bigr\rangle=0.$ (50)
@@
$F_{ij}^{\textrm{mod},ca}(t)=\bigl\langle[f_{i}(t),f_{j}(0)]_{+}[f_{i}(t),f_{j}(0)]_{+}^{\dagger}\bigr\rangle$
$=\bigl\langle[f_{i}(t),f_{j}(0)]_{+}[f_{i}^{\dagger}(t),f_{j}^{\dagger}(0)]_{+}\bigr\rangle$
$=(1-\delta_{ij})2\bigl\langle P_{i}^{-}P_{j}^{-}\bigr\rangle$
$+(1-\delta_{ij})2\operatorname{Re}\bigl\langle\mathcal{U}_{i}(t,0)f_{i}f_{j}f_{i}^{\dagger}\mathcal{U}_{i}^{\dagger}(t,0)f_{j}^{\dagger}\bigr\rangle=0.$ (51)
@@ V Square of the anticommutator
$F_{ij}^{\textrm{sq}}(t)=\bigl\langle\bigl([f_{i}(t),f_{j}^{\dagger}(0)]_{+}\bigr)^{2}\bigr\rangle$
$=\bigl\langle[f_{i}(t),f_{j}^{\dagger}(0)]_{+}[f_{i}(t),f_{j}^{\dagger}(0)]_{+}\bigr\rangle$
$=\bigl\langle f_{i}(t)f_{j}^{\dagger}(0)f_{i}(t)f_{j}^{\dagger}(0)\bigr\rangle+\bigl\langle f_{j}^{\dagger}(0)f_{i}(t)f_{j}^{\dagger}(0)f_{i}(t)\bigr\rangle$
$+\bigl\langle f_{i}(t)f_{j}^{\dagger}(0)f_{j}^{\dagger}(0)f_{i}(t)\bigr\rangle+\bigl\langle f_{j}^{\dagger}(0)f_{i}(t)f_{i}(t)f_{j}^{\dagger}(0)\bigr\rangle.$ (52)
@@
$F_{ij}^{\textrm{sq}}(t)$ $=\bigl\langle\mathcal{U}_{i}(t,0)f_{i}f_{j}^{\dagger}\mathcal{U}_{i}(t,0)f_{i}f_{j}^{\dagger}\bigr\rangle$
$+\bigl\langle\tilde{\mathcal{U}}_{i}(t,0)f_{i}^{\dagger}f_{j}\tilde{\mathcal{U}}_{i}(t,0)f_{i}^{\dagger}f_{j}\bigr\rangle^{*}.$ (53)
@@
$F_{ii}^{\textrm{sq}}(t)$ $=\bigl\langle\mathcal{U}_{i}(t,0)\mathcal{U}_{i}(t,0)P_{i}^{-}\bigr\rangle$
$+\bigl\langle\tilde{\mathcal{U}}_{i}(t,0)\tilde{\mathcal{U}}_{i}(t,0)P_{i}^{+}\bigr\rangle^{*}.$ (55)
@@ VI Results and Discussion
At half filling $n_{d}=n_{f}=0.5$, the contributions $F_{0}(t)$ and $F_{1}(t)$ in (58) are equal and the total OTOC $F(t)$ is real. But, the partial contributions from the sites with empty $f$ particle states, $F_{0}(t)$, and with occupied $f$ particle states, $F_{1}(t)$, display intriguing time dependencies. For $U<0.44$, the real part of both characteristic functions (72) is positive at frequencies where the imaginary part vanishes and the winding number is zero (see Appendix B). The OTOC function (57) displays exponential decay at large times for high temperatures, which is replaced by a power law as $T\to 0$ (Fig. 4). The corresponding logarithmic derivative
@@
approaches a constant value at large times for high temperatures, with a decay rate within the “chaos bound” $|L_{0}(t)|/(2\pi T)<1$  [^1] Despite the chaos bound of [8, 9] is applicable for the short timescale, we compare the decay rate with this temperature bound to elucidate thermal effects in relaxation., whereas it displays $1/t$ dependence for low temperatures, which confirms the power law asymptotics of the OTOC. Consequently, the corresponding “density-of-states”  [^2] Now, DOS is defined by the real part of OTOC because (52) omits factor $\mathrm{i}$” in contrast to the one-particle Green’s function (14)
$A(\omega)=\frac{1}{\pi}\operatorname{Re}\int_{0}^{+\infty}\mathrm{d}t\;\mathrm{e}^{\mathrm{i}\omega t}F_{ii}^{\textrm{sq}}(t)$ (74)
@@
Figure 4: (a) OTOC $F(t)$, (b,c) normalized logarithmic derivative $L_{0}(t)/(2\pi T)$, and (d) the corresponding “density-of-states” $A(\omega)$ for $U=0.4$, $n_{f}=0.5$, $n_{d}=0.5$.
@@
where $W=\Delta t\to 0$ is the time discretization, and the long-time asymptotic becomes more complicated (Fig. 5) with oscillations of the logarithmic derivatives $L_{0}(t)$ at high temperatures similar to the one for the single-particle propagators for $U>0.866$ (Fig. 2). With decreasing temperature, these oscillations are replaced by $1/t$ asymptotics, because we are still in the metallic phase $U<\sqrt{2}$. Now the OTOC’s “density-of-states” displays a two peak structure separated by a pseudogap at high temperatures with the development of the power law singularity in the center of the pseudogap as $T\to 0$. The sharp peak in the “density-of-states” is observed at low temperatures up to $U$ values where the “density-of-states” becomes negative, which reflects an increase in the quantum effects for these cases (such as Kirkwood-Dirac quasiprobabilities (KDQ) [40, 41, 42], which can also become negative).
For $U>0.866$, the imaginary part of the characteristic function (72) crosses the zero line at frequencies where $1\mp U\operatorname{Re}\mathcal{G}_{0,1}^{r}(\omega)=0$, but for $U<1.235$ the winding does not change because the real part is still positive at these points, whereas for $U>1.235$ the real part becomes negative and the winding number is equal to 2-2 (see Appendix B) resulting in a more complicated long time behavior of the OTOCs and their logarithmic derivatives with damped oscillations (Fig. 6). Now, the OTOC’s DOS is more robust, displays weak temperature dependence, and possesses negative values in a wide frequency interval. For large values of interaction $U$, the quantum correlations become dominant and the “flip” of an $f$-state produces strong disturbance for the itinerant $d$-electrons, which can subsequently lead to local heat fluctuations. Such quantum heat fluctuations can be associated with negativity of quasiprobability distribution, e.g. Kirkwood-Dirac quasiprobability, which, on the other hand, can be a signature of suppressing of the decoherence and of quantum chaos and scrambling [43, 44].
However, the sum rule (integral over frequency) for OTOC’s DOS is conserved and is equal to one:
$\int_{-\infty}^{+\infty}\mathrm{d}\,\omega A(\omega)=F_{ii}^{\textrm{sq}}(0)=1,$ (76)
which means that the appearance of negative DOS at small frequency values leads to a DOS enhancement elsewhere. For the first moment we obtain
$\int_{-\infty}^{+\infty}\mathrm{d}\,\omega\,\omega A(\omega)=\mathrm{i}\frac{\mathrm{d}F_{ii}^{\textrm{sq}}(t)}{\mathrm{d}t}\bigg|_{t=0}=2(Un_{d}+E_{f}-\mu),$ (77)
which equals zero at half filling $n_{d}=n_{f}=0.5$, and one can proceed similarly and systematically to determine higher-order moments.
@@
It should be noted, that the Szegő-Widom theorem for block Toeplitz matrices [37] is applicable for smooth functions. For this case in the fermionic FKM, the functions are continuous at high temperatures. At low temperatures, both the matrix elements in (68) and characteristic functions (72) develop step-like features from the Fermi factors and the OTOC’s long time behavior is now governed by localization singularities given by the Anderson orthogonality catastrophe [45, 35, 34]. These cases of three different winding numbers for the characteristic function are also distinguished by the behavior of the logarithmic derivative at low temperatures (see Appendix C, Fig. A.6).
At high temperatures, the oscillatory behavior of the logarithmic derivative is similar to the one seen in [12, 13] for quantum spin models behind the light cone after the wave front passes. This analogy is plausible, because we consider a local OTOC function, which is always behind the wave front and at high temperature the effects of the Fermi statistics become negligible and the electrons behave like Pauli particles. Nevertheless, the localized $f$ particles act as impurity scattering centers for the itinerant $d$ electrons and the destruction of the oscillations of logarithmic derivative for large Coulomb interaction $U$ is similar to the one observed in [13] with the an increase of disorder, when one observes the post wavefront ripples damped by disorder. This static disorder can protect the fast scrambling, which can be thought of as a breaking of the “bound on chaos”, when at long times we have $|L_{0}(t)|>2\pi k_{B}T/\hbar$, which, is also manifested by the negative OTOC “density-of-states” $A(\omega)$, see Fig. 6. According to Yunger Halpern et al. [41], OTOCs can be considered as moments of the Kirkwood-Dirac quasiprobabilities which implies that negative values of the “density-of-states” $A(\omega)$ are signatures of quantum information scrambling.